I am a 3-digit number. My ones is 2 more than my hundreds and my tens is 3 less than my ones.The sum of my digits is 13.What number am I?
Answer by lwsshak3(11628)

I am a 3-digit number. My ones is 2 more than my hundreds and my tens is 3 less than my ones.The sum of my digits is 13.What number am I?
**
h=hudreds digit=h
u=ones digit=h+2
t=tens digit=u-3=h+2-3=h-1
..
h+u+t=13
h+(h+2)+(h-1)=13
3h+1=13
3h=12
h=4
u=h+2=6
t=h-1=3
number: 436